Prev Next

SW&IT Meeting Minutes 2018-12-17

Attendees: Andrew, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every, Do not edit this table, use entries in smartsheets. 20181119
LR20181119 - Bec, Matthieu 64.0
LR20181119 - Colson, Andrew 78.0
LR20181119 - Edmeades, David 64.0
LR20181119 - Sola, Igor 53.0
LR20181119 - Kubanek, Petr 64.0
LR20181119 - Hooper, Stephen 64.0
LR20181119 - Zhang, Yang 40.0
LR20181119 - Zhang, Xianyu 0.0

20181203
LR20181203 - Bec, Matthieu 65.0
LR20181203 - Colson, Andrew 48.0
LR20181203 - Edmeades, David 24.0
LR20181203 - Sola, Igor 24.0
LR20181203 - Kubanek, Petr 72.0
LR20181203 - Hooper, Stephen 80.0
LR20181203 - Zhang, Yang 80.0
LR20181203 - Zhang, Xianyu 0.0

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input from secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.


Application(s) Primary Secondary Week
AzCam
AC (2)
Petr (2)
2
Weatherstation
AC (2)
Petr (2)
2
FACSUM,WEATHERSTATION
AC (2)
Matthieu (2)
4
ALH Igor(2)
Matthieu (2)
1
INDI Igor (2)
Matthieu (2)
3
Status Server Stephen (2)
Matthieu (2)
2
MODS
Yang (3)
Igor (1/2) 2
Telem-VIZ
Yang (1)
Stephen (1) 3
particle monitors AC Petr 4

News from management

  • PEPSI/pol high profile run in January: AGW3 & 4 must be ready
  • ARGOS half-night Dec 22 - confirmed
  • Time Sheets for Pay Period End Date December 30, 2018 must be submitted and approved no later than 5:00 pm on Tuesday, December 18, 201

Previous Action Items

  • Skill set list (DE)
  • UA time (AC)

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • seeing monitor upgrade
  • end of year support
  • network dropout - antenna monitoring?
  • iLocater Single-Mode Data Fibers - CRA/CRB to Electronics Shop
  • BCU f/w upgrade documentation
  • Dust monitor status
  • AO improvement: Luci kwds
  • AO cabinet switch upgrade - follow up
  • wifi summit upgrade (note: end-of-year deadline)
  • Tucson CR power overhaul plan (note: end-of-year deadline)
  • n507/obs upgrade (note: end-of-year deadline)
  • synology new drives
  • PGray mirror cover req. feedback
  • Installing python packages on flao-sxwfs
  • support aligned with mountain schedule (starting Tuesday instead of Monday)? (PK)

This week in github

Team reports

PK

  • OAC pair programming (with AC), fixing socket communication for the new housekeeping
  • IT7319 - leap second file
  • PM boxes - learning VxWorks details, test mode, laptop VxWorks booting, testing serial connection (which so far failed to identify problem). We plan to test the VxWorks on mountain when time will be available (need 2-4 hours to see what's wrong, probably after New Year). Details TBC
  • Online C/C++
  • GCS crash fix (inproper ue of asctime and localtime in multithread environment, combined with creative programming to remove trailing line feed in ctime instead of calling strftime with %c). next:
  • GCS
  • AOC
  • support VxWorks tests

AC

  • OAC svn to git complete
  • v5.1 now running in lab
  • AGW4 going to the mountain on 12/18; AGW3 TBD
  • Broken relay in controller board replaced
  • Particle monitor telemetry

YZ

  • Updated the anonymous account google doc with LBC and MODS login.
  • Working on the AO-Cravat web interface design. Made some more progress.
  • Working on MODS wiki overview documents, a tech talk to be scheduled on either Thursday or Friday (Dec 20th or 21st).

MB

  • OC, WPC, management
  • iLocater FDR
next:
  • catching up

XZ

last week:
  • Monday, Tuesday vacation, but worked some hours with OVMS+TT, ASM SW related to short dust check and zoom calls
  • Power Backplane plan discussion and write simple test plan
  • Document the old crate BCU memory map
  • 4D interfermeter SW interface
  • 4D HW & SW requirements dicussion
  • Current 4D resolution check (1 frame in x)
  • Fastlink fiber swaps configration and debug
This week:
  • countiue document the old crate BCU memory map
  • analysis more measurement for current 4D resolution check
  • OVMS+TT meeting?
  • AOS IT?
  • AO support

DE

  • Fill-in CDW rep has realized situation re: drives and is pushing for overnight shipping
    • Will endeavor to fit future purchases under the Pcard limit and order from Amazon. This was ridiculous.
    • Should have them this week, will install asap
  • Rm507-2 UPS died, swapped it with spare unit
    • Doing some tests on it to see if it can be salvaged
  • Autocad policy hammering out; Jonathan is angling for something
    • I believe that I can do everything Jonathan thinks needs to be done re: paths manually; few enough installs that it is not overly burdensome
    • Other users seem happy with my 2018 installs
  • Laura is functional by bypassing dock's USB for kb/mouse.
    • I have some tasks from Dell, but she hasn't had time to give me the laptop
  • Ordered replacement batteries for server room UPS

SH

  • continuing mountain network woes.
  • Ordered monitor for Bucki
  • AO cabinet switch configuration (try to feed 100Mbps in the lab) -- received switch Thursday.
  • Microwave information collection -- webcam
  • DIMM computer replacement -- quotes in, discussed, waiting.
  • ARGOS computer spare -- quotes in, discussed, waiting.
  • Loop problems and CRB switch issue -- luci wiring, but no mountain access for a couple of days.
  • mountain VM for SOUL telemetry.
  • meeting with SMT about WiFi.

next:
  • LBTI computer IP addresses -- check back with Paul (on NTP time as well).
  • refine info page -- add info, better history, less bandwidth, mobile site refinement -- not very important so lower priority.
  • SOUL work for DX -- schedule.
  • AO cabinet switch work -- schedule.

ISS

  • Last Week
    • Finish AFC Debugging tool
    • Luci Joiner for Argos/ESM scripts (ongoing)
  • This Week
    • Put the AFC debugging in production and have Dave T. use it.
    • Look at RTD issues #7473, #7488 and #7512.
      • Unable to use AFC Align Pinholes to Align Field Stop For N30 Camera
      • Slow subtraction in RTD
      • LUCI-2 skips field image during MOS acquisition
    • Luci joiner for Argos/ESM scripts
Topic revision: r17 - 21 Dec 2018, PetrKubanek
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback